[Brahms-l] Shift report 20050623 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Jun 23 2005 - 07:58:56 EDT
Shifter leader was AG
Shifters were 


1. Beam:
   - Physics Beam for ~4.5 hours, Fill pattern: 111x111_p3
   - Polarization at beginning (02.50) B/Y: -43.32 +/- 2.37, -48.64 +/- 2.21
   - New polarization (05.55): B/Y: -46.66 +/- 2.40, -52.15 +/- 2.33

2. Data: 
   - Runs 15759 - 15766
   - Trig2(723k),Trig3(162k),Trig4(113k),Trig5(778k)
   - Pedestal run done and installed in tpcped, Non pedestal run done.

3. Problems:
   - In TOFW and CALFB the actual voltage is slightly more negative than the demanded one.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jun 23 07:59:10 2005

This archive was generated by hypermail 2.1.8 : Thu Jun 23 2005 - 07:59:23 EDT